OpenStudy (anonymous):

If 854 = 12, 467 = 38, then 683 =?

Parth (parthkohli):

854 is exactly 2 times 467. 467 is 38. 683 is the mean of the numbers. So, mean of 12 and 38 will be the answer.

Parth (parthkohli):

Mean of 12 and 38 is 25.

OpenStudy (anonymous):

But then what 854=12 means

Parth (parthkohli):

It's a pattern.

OpenStudy (lgbasallote):

why is 467 = 38 then?

Parth (parthkohli):

Wait wait wait... 683 is not the mean..

Parth (parthkohli):

And, it's not 2 times...

Parth (parthkohli):

I dunno this time.

OpenStudy (anonymous):

18 ?

OpenStudy (anonymous):

@FoolForMath how did u get that please please please tell me

Parth (parthkohli):

It's 10..

Parth (parthkohli):

Option c.

OpenStudy (anonymous):

What does option c mean?

OpenStudy (lgbasallote):

this reminds me of lgbariddles

OpenStudy (anonymous):

@lgbasallote what does that mean

OpenStudy (anonymous):

@.sam need help

OpenStudy (anonymous):

@.Sam. need help

Parth (parthkohli):

It's some kinda riddle.

Parth (parthkohli):

It's not mathematical, that's what I think

OpenStudy (lgbasallote):

854 = 12 8 x 4 - (5 x 4) 467 = 38 7 x 4 + (6 + 4) lol weird pattern doesnt make sense though

OpenStudy (anonymous):

@.Sam do you have any opinions

Parth (parthkohli):

I found the answer by googling. It's 10.

Parth (parthkohli):

I don't understand why..

OpenStudy (anonymous):

@Diyadiya can u sort it out?

OpenStudy (anonymous):

Never mind @ParthKohli

OpenStudy (lgbasallote):

lol haha that makes sense

OpenStudy (anonymous):

Please .... someone help...

OpenStudy (lgbasallote):

haha got it :P

OpenStudy (anonymous):

OK forget .... if anyone gets any clues please message me ....

Parth (parthkohli):

No, when I googled it was 10.

OpenStudy (lgbasallote):

should be 18...only pattern i can see

Parth (parthkohli):

Can you post it again lgba?

OpenStudy (diyadiya):

Ohhhhh!!!!!!!!!!!! Yeah 18

OpenStudy (anonymous):

come o tell me the pattern lol

Parth (parthkohli):

Yes, I think he is right..

OpenStudy (lgbasallote):

854 = 12 5 x 4 - 8 467 = 38 7 x 6 -4 683 = ? 8 x 3 - 6 = 18

Parth (parthkohli):

5 * 4 - 8 = 12 8 * 3 - 6 = 18

OpenStudy (anonymous):

So easy ;)

OpenStudy (lgbasallote):

that was hard fool :P

OpenStudy (anonymous):

great work @lgbasallote thank you very very much

OpenStudy (anonymous):

@FoolForMath then why did'nt u reply

OpenStudy (lgbasallote):

yup googled it...18 is right :D